If the coefficients of the rth and (r+1)th term in the expansion of (3+7x) raised to power 29 are equal, then r=..?

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

I will give u a hint here:

Expansion (x+y)n :

(i)        There are (n+1) terms in the expansion.

(ii)        In each term, sum of the indices of ‘x’ and ‘y’ is equal to ‘n’.

(iii)        In any term, the lower suffix of ‘c’ is equal to the index of ‘y’, and the index of x = n-(lower suffix of c).

(iv)             Because  nCr = nCn-r,

    so we have :

nCo = nCn

nC1=nCn-1

nC2=nCn-2 etc.

It follows that the coefficients of terms equidistant from the beginning and the ends are equal.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com